OpenStudy (anonymous):

LeAnn has 20 coins in her piggy bank. All the coins are either pennies or dimes. The total value of the coins is 83¢. How many pennies are there? How many dimes?

OpenStudy (anonymous):

@primeralph

OpenStudy (primeralph):

I don't even know what a dime and penny hold in terms of monetary value. I usually throw them away.

OpenStudy (anonymous):

-_-

OpenStudy (primeralph):

@e.cociuba

OpenStudy (anonymous):

dimes= 10 cent pennies=1 cent

OpenStudy (primeralph):

@amistre64 @satellite73 Anybody know what a penny is?

OpenStudy (anonymous):

1 cent!

OpenStudy (amistre64):

i have tons of them in the floorboard of my car :)

OpenStudy (anonymous):

Can someone help me ? -____-

OpenStudy (amistre64):

p+d = 20 1p + 10d = 83 solve the system

OpenStudy (primeralph):

A penny is 1 cent? What's that good for?

OpenStudy (anonymous):

@arilove1d

OpenStudy (amistre64):

i jsut gave you the equations you need to relate :/

OpenStudy (anonymous):

OMG i just did this

OpenStudy (anonymous):

o.o

OpenStudy (amistre64):

p+d = 20 1p + 10d = 83 solve the system

OpenStudy (anonymous):

Well sorry , you DONT Have to help me anymore then , since you wanna catch an attitude.

OpenStudy (amistre64):

let p=10+n let d=10-n

OpenStudy (anonymous):

10 dimes and 3 pennies ? .-.

OpenStudy (anonymous):

whoa i was just saying i just learned it

OpenStudy (amistre64):

(10+n)+(10-n) = 20 ; this is true for any value of n (10+n) + 10(10-n) = 83 ; find n

OpenStudy (anonymous):

Oh... sorry..

OpenStudy (amistre64):

10+3 does not equal 20 coins

OpenStudy (anonymous):

its ok i should have rephrased that

OpenStudy (anonymous):

I know , I mean 8 dimes and 3 pennies because 8 dimes is 80 cent + 3 pennies = 83 cent :D

OpenStudy (amistre64):

there are many combinations that work for one but not the other .... guessing could take all day (10+n) + 10(10-n) = 83 ; find n

OpenStudy (anonymous):

Hmmm... 10 ?

OpenStudy (amistre64):

good luck ...

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

p+d = 20 1p + 10d = 83 subtract 1st equation from second: p-p + 10d-d = 83-20 9d = 63 d = 7 p = 20-d = 13 LeAnn has 7 dimes and 13 pennies

OpenStudy (primeralph):

I don't care, this person is broke. Really.

OpenStudy (anonymous):

Oh im late .......

OpenStudy (anonymous):

Lol , who ?

OpenStudy (primeralph):

LeAnn.

OpenStudy (anonymous):

BOL!

OpenStudy (anonymous):

.-.

OpenStudy (anonymous):

I didnt even see that what ari put... THANK YOU ARIII! :)

OpenStudy (anonymous):

YOUR WELCOME !!!

OpenStudy (anonymous):

>:D